Amy's brother is half the age of Amy. Amy's mom is twice the age of Amy and seven years older. The sum of their age is 56. What is Amy's age? What equation would you make for this?
Answer by ankor@dixie-net.com(22740)

Let a = Amy's age
Let b = Brother's age
Let m = Mom's age
:
Write an equation for each statement:
:
Amy's brother is half the age of Amy.
b = .5a
:
Amy's mom is twice the age of Amy and seven years older.
m = 2a + 7
;
The sum of their age is 56.
a + b + m = 56
Replace b with .5a, replace m with (2a+7)
a + .5a + (2a+7) = 56
Combine like terms, subtract 7 from both sides
3.5a = 56 - 7
3.5a = 49
a = 49/3.5
a = 14 yrs is Amy's age
:
:
You can check this for yourself, find b and m and find the total of all three ages